OpenStudy (anonymous):

34/6= 2x+1/3 i really need help thanks everyone!!!!!!

OpenStudy (anonymous):

so first off isolate your x by subtracting 1/3 from both sides

hero (hero):

1. Reduce 34/6 to 17/3 17/3 = 2x + 1/3 2. Subtract 1/3 from both sides: 17/3 - 1/3 = 2x 16/3 = 2x 3. Finish solving from there.